Jackson Inc. produces leather handbags. The production budget for the next four months is: July 5,000 units, August 7,000, September 7,500, October 8,000. Each handbag requires 0.5 square meters of leather. Jackson Inc.'s leather inventory policy is 30% of next month's production needs. On July 1 leather inventory was expected to be 1,000 square meters. Leather is expected to cost $5 per square meter in June, but go up to $6 per square meter in July. What is the expected cost of leather purchases in July?
Answer #1

Solution:

July Production Needs of Leather = 5000 units *0.5 = 2500 Square meters

July 1 Leather inventory = 1000 square meters

July 31 Leather Inventory = August Product needs * 30% = 7000*0.5 *30% = 1050 square meters

Leather purchases in July = July 31 Leather inventory + July Production Needs of Leather - July 1 Leather inventory

= 1050 + 2500 - 1000 = 2550 square meters

expected cost of leather purchases in July = 2550 * $6 = $15,300